OpenStudy (kj4uts):

Suppose a rock is thrown upwards from the top of a building 200 feet high with an initial velocity of 96 feet per second. The height (in feet) of the rock above the ground t seconds after being released is given by the formula h(t) = -16t^2 + 96t + 200 1.When will it hit the ground? 2.How high will it be after 2.1 seconds? 3.What is the maximum height?

OpenStudy (anonymous):

1. It will hit the ground when h(t) = 0 so you need to solve \[-16t^2+96t+200=0\]

OpenStudy (anonymous):

try the quadratic formula

OpenStudy (kj4uts):

Would it be t=3-√ 43/2=-1.63 and 3+√ 43/2=7.63

OpenStudy (anonymous):

yes, and since time can't be negative, it hits the ground at 7.63 seconds.

OpenStudy (anonymous):

For 2, plug in 2.1 for t to find h(2.1).

OpenStudy (kj4uts):

331.04

OpenStudy (anonymous):

yes. and then for 3 there are two steps. Use \(t=-\frac{ b }{ 2a }\) to find the time of max height. Then plug that time into the h(t) equation

OpenStudy (kj4uts):

Im a little confused as to what I plug into b and a?

OpenStudy (anonymous):

the same values you used in the quadratic equation a = -16, b = 96, c = 200

OpenStudy (anonymous):

\[t=-\frac{ 96 }{ 2(-16) }\]

OpenStudy (kj4uts):

3

OpenStudy (anonymous):

now plug 3 in for t in the original equation

OpenStudy (kj4uts):

Ok I got 344 thank you for explaining this step by step :)

OpenStudy (anonymous):

yeah that's it. you're welcome. :)